OpenStudy (waheguru):

How Do I find the slope of the given points whthout making a graph c(14,8) m(21,22)

OpenStudy (saifoo.khan):

Use the slope formula.

OpenStudy (waheguru):

rise/run = slope?

OpenStudy (saifoo.khan):

You got that right bro. ;)

OpenStudy (waheguru):

So 22+8/ 14+21 = 30/35 =6/7 ? But this is the wrong answer @saifoo.khan

OpenStudy (asnaseer):

you need to subtract to work out the rise and run

OpenStudy (waheguru):

Oh, We need to substact?

OpenStudy (saifoo.khan):

This should help: http://ceemrr.com/Geometry1/EquationsOfLines/image014.gif

OpenStudy (asnaseer):